OpenStudy (anonymous):

Question below:

OpenStudy (anonymous):

If \(\cos 2 \alpha=\dfrac{3\cos2 \beta-1}{3\cos 2 \beta}\) then \(\tan \alpha=?\)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

use the identity (tan x)^2 = (1-cos2x)/(1+cos2x) and then proceed...

OpenStudy (anonymous):

Let me see.

OpenStudy (anonymous):

I got \[\dfrac{1}{6 \cos 2\beta-1}\]

OpenStudy (anonymous):

check your simplification

OpenStudy (anonymous):

I am still getting the same answer.

OpenStudy (anonymous):

Of course I will have to take square root to get tan alpha

OpenStudy (anonymous):

@matricked ?

OpenStudy (anonymous):

you can further use thge identity cos2y = (1- (tany)^2) / (1- (tany)^2) and proceed further...

OpenStudy (anonymous):

So my simplification is right?

OpenStudy (anonymous):

it seems...

OpenStudy (anonymous):

Further simplification gives me: \[\dfrac{\sec^2 \beta}{7-5\tan^2 \beta}\]

OpenStudy (anonymous):

just take the root then ..

OpenStudy (anonymous):

by the way can you recheck the question is somewhere 5 instead of a 3

OpenStudy (anonymous):

\[a)\sqrt{2} \tan \beta \;\;b)\tan \beta \;\;c)\sin 2\beta \;\;d)\sqrt{2}\cot 2\beta\]

OpenStudy (anonymous):

The question is right.

OpenStudy (anonymous):

Shall i try my hand at solving this?

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

sure help him...

OpenStudy (anonymous):

Is any information given about Angles alpha and beta

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Like they are acute or less than 90 greater than 90 ,acute (like these)

OpenStudy (anonymous):

It must be given otherwise u can't proceed ahead!

OpenStudy (anonymous):

??

OpenStudy (anonymous):

not given

OpenStudy (anonymous):

Check the heading of the question it must be or we have to assume things here

OpenStudy (sidsiddhartha):

then assume alpha beta as acute angle

OpenStudy (anonymous):

@OptimusPrime_ can you check the question once again...

OpenStudy (anonymous):

it is right.

OpenStudy (anonymous):

not about the question if any extra information is given check it seriously @sidsiddhartha negative or positve acute angle

OpenStudy (anonymous):

can you reveal the correct options...

OpenStudy (sidsiddhartha):

its sqrt(2)

OpenStudy (anonymous):

I will assume it is positive acute angles for answer sake!

OpenStudy (sidsiddhartha):

if they are acute

OpenStudy (anonymous):

No idea..qn is from a test I took today. I couldnt solve it during test.

OpenStudy (anonymous):

I will help you! wait

OpenStudy (anonymous):

If A and B are positive angles ( that's what i am assuming they must lie in the first Quadrant because theta is less than 90) From this we can interpret that sin A, cos A, tan A, tan B > 0

OpenStudy (anonymous):

@OptimusPrime_ if you don't mind you are studying in which board

OpenStudy (anonymous):

Sorry how is that relevant here?

OpenStudy (anonymous):

\[\cos 2\alpha = 1 - 2\sin ^{2}\alpha \]

OpenStudy (anonymous):

Are u familiar with that identity?

OpenStudy (anonymous):

Yes. Proceed.

OpenStudy (anonymous):